Rena update (update 145)

6 January 2012, 10:30am

A total of 17 containers were removed from the no. 5 and 6 holds on Rena yesterday.

Salvage

  • Salvors plan to continue with container removal operations today while the weather remains favourable. This brings the total estimated number of containers still on board to 881.
  • 389 containers in total have been removed from Rena, with 98 in total believed to have been lost overboard.
  • Salvors are preparing the ship for severe bad weather this weekend, with a low pressure system now expected to arrive on Saturday evening, carrying through to Sunday, with the worst weather now expected on Monday (9 Jan). This is forecast to result in a maximum wave height of 6-7m, which is among the worst weather to be forecast so far.
  • In expectation of the bad weather front, salvors will tomorrow (Saturday) bring the crane barge Smit Borneo back into port. Salvors will use the time to complete routine maintenance on the barge and carry out other work. The support vessel Go Canopus will remain on site to monitor the condition of Rena.
  • Containers on the bow of the vessel continue to remain lashed down and those that are safely accessible have been fitted with transponders in anticipation of the severe weather front coming through.
  • Work is continuing on erecting a temporary gangway between the forward and aft sections of ship, with a more permanent structure to be built.
  • High gas levels detected in the no.5 hold late yesterday caused work to be stopped for safety reasons. Breathing apparatus and other equipment is on the vessel to help manage the hazard, and levels are continuing to be monitored today.
  • No diving work took place yesterday due to the conditions, but divers hope to be able to inspect the recent damage to the vessel today. This will however be dependent on the conditions and will only occur when it is safe for them to enter the water.
  • The ship remains in a fragile but stable condition, with electronic sensors on board showing no significant change.
  • Although effectively in two parts, the forward and aft sections of the vessel remain firmly grounded on the reef.
  • A 3NM exclusion zone remains in place around the vessel, as does a 1500 foot vertical air exclusion zone. Boaties are advised that the zone may need to be extended if the forecast bad weather results in more containers or debris being lost from Rena, as this will pose a hazard to shipping. If this occurs, navigational warnings will also be issued via maritime radio.

Containers

  • Braemar Howells are preparing vessels and other equipment in readiness to respond to potential container and debris loss from the Rena with the forecast bad weather. This includes preparing tow wires, ropes and trawl nets to collect debris and containers that may come off the vessel.
  • Work on collecting container debris from Motiti Island is continuing today, comprising mainly timber, milk powder and plastic. Teams are also collecting debris from other beaches.
  • Two teams yesterday found debris at Papamoa Beach (mainly milk powder) and this is being collected. The public are again urged to please not touch the debris, as it makes recovery more difficult and may pose a hazard. Instead, anyone finding container debris is asked to please report it via the oil spill number: 0800 645 774.
  • Timber debris reported by fishermen in the water near Motiti Island yesterday and scattered over a wide area has been collected. Only small debris remains.
  • No sonar work was carried out yesterday due to technical problems, however sonar work has resumed today.

Oil spill response

  • Responders remain on standby and are ready to ramp up in anticipation of possible further releases of oil from Rena following the forecast bad weather this weekend.
  • A metallic sheen of oil was observed on both sides of the ship during this morning’s over flight, measuring about 50m wide and extending for about 4.2nm towards the north of the vessel.
  • Work is continuing at Leisure Island today and preparations to resume work at Matakana Island next week are continuing.
  • Any sightings of oil can be reported to 0800 645 774, as lines will continue to be monitored over the weekend. Similarly, any reports of oiled wildlife should be reported to 0800 333 771.

All media releases

Subscribe Follow us News feed

Call the MNZ media line

Media centre

Additional information for media.
Media centre

Media archive

View our historic media releases.

Media archive